# An electron-hadron collider at the high-luminosity LHC

### v2

@@ 1 Introduction
The Large Hadron-electron Collider (LHeC) [1, 2] will open new frontiers in high-energy physics as it will become a unique laboratory for studies of deep-inelastic scattering (DIS) and Higgs boson, electroweak, top quark and beyond the Standard Model (BSM) phenomena in a vast range of virtualities $Q^{2}$, from very close to zero up to $10^{6}$ GeV2, and for a much extended range of $x$ – the fractional longitudinal momentum of partons. Recently, an intense electron beam from a novel energy recovery linac (ERL) was proposed to collide with a proton or ion beam from the circular, 27-km circumference High-Luminosity Large Hadron Collider (HL-LHC) at CERN, in the redesigned P2 interaction region capable of hosting electron-hadron or hadron-hadron collisions, concurrently with hadron-hadron collisions in other HL-LHC experiments [3].
In the Conceptual Design Report (CDR) of the LHeC, it is assumed that a three-pass ERL delivers a 50 GeV high-current electron beam with an intensity of up to 20 mA [2], and in the most recent update of the LHeC proposal, its commissioning is proposed several years after the completion of the HL-LHC programme [4]. In this paper, we propose the use of a 20 GeV single-pass ERL with yet a higher beam current of 60 mA, and with a circumference of a third of the LHC’s. This choice of design is motivated by the availability of the technology to operate a single-pass ERL (although presently at lower beam currents) and by the overall simplification, in particular, in the design of the adequate machine-detector interface owing to much lower levels of synchrotron radiation, allowing much earlier commissioning of the high energy electron-hadron collisions at CERN. The electron beam will be accelerated to a full energy of 20 GeV by one pass in two 10 GeV linacs (see Fig. 1), and after collision with the clockwise circulating LHC proton beam, it will be guided back to the linac structures in counter-phase to be completely decelerated within the energy recovery process.
@@
The technical feasibility of such a single-pass design has been demonstrated in a number of test facilities and linacs that operate in the range of several mA beam current (see, e.g., [5]). Fig. 2 summarises ERLs that are already operational, in the phase of commissioning or planned.
Figure 2: Different ERL projects worldwide that are already operational or in their commissioning / planning phase. The red and green markers indicate the baseline LHeC design and the ERL configuration described in this paper, respectively. Figure adapted from [5].
Note that the cost of building this ”phase-one” LHeC as well as its running costs are significantly lower than for the final LHeC configuration. Such a staged realisation of the LHeC project provides a huge benefit not only in reducing financial (and personnel) stress but first of all in the availability of concurrent electron-hadron collisions already during Run5 of the current HL-LHC programme. Consequently, apart from the possibility of carrying out a wide range of unique and important studies almost 10 years earlier, the results of this electron-hadron experiment could also be used to improve many important measurements in other (hadron-hadron) experiments at the HL-LHC in a timely manner.
2 The LHeC interaction region
@@
Schematically, the concept is shown in Fig. 1 – the geometry of the ERL will remain the same: two superconducting linacs are connected by return arcs on both sides – details about the general layout can be found in [6] and [7]. However, for the phase-one case with reduced electron energy several parts needed to reach higher energies are left out: arcs from 3 to 6 will not be installed, nor will the so-called bypass needed in the final 50 GeV design to guide the beam outside the particle detector, which will only be installed in a later staging step. Beam spreaders and re-combiner modules will also not be present in the phase-one electron beam lattice.
On the other hand, a number of key elements will define the geometry of the ERL from phase one and will remain untouched for all energy steps. The linacs, the interaction region (IR) and the beam separation scheme will be designed to be valid for any operation energy. As described in detail in [6] the proton and electron beams are separated after the interaction point through their different beam rigidities. A combination of a weak dipole field embedded inside the detector and the combined function effect of the off-centred mini-beta quadrupoles of the electron lattice acts as an effective separation scheme. The actual value of the dipole strength and the offset of the quadrupoles were optimised for the lowest synchrotron radiation and thus equal effective bending radii in the three magnets [8], and for the phase-one scenario described here, the electron bending radius of this separation scheme is of about 1500 m.
@@
Figure 3: The interaction region of electron and proton beams: in black the mini-beta focusing scheme of the electrons is shown, including the detector dipole field that is part of the beam separation. Both are embedded within the free space of the HL-LHC proton inner triplet lattice, marked in blue and red. Figure adapted from [9].
The interaction region of the $ep$ collisions is shown schematically in Fig. 3. Given the comparatively low energy of the electrons, a gradient of approximately g=20g=20~T/m and a length of 1.8 m are sufficient to meet the optics requirements of the electron quadrupoles. These magnets can therefore be made very compact, and the dimensions of the electron mini-beta insertions have been chosen so that they will fit easily in the free space of the IR. Only the first quadrupole of the proton lattice, “Q1” in the figure, will have to be replaced by a new design that allows a field free region for the electrons as their separation is not large enough yet to be guided completely outside the cryostat. For details of the technical design of this new magnet we refer to [2]. With the separation scheme explained here, head-on collisions between the proton and electron beam are foreseen, to avoid luminosity reduction because of the geometric loss factor from an eventual crossing angle.
@@
Figure 4: Local distortion of the proton optics due to the influence of the electron quadrupoles before (red) and after a local compensation of the effect (blue) for the case of a 20 GeV electron energy. Figure extracted from [9].
Beyond the optics distortion from the electron gradient fields that have a certain influence on the proton optics, the effect of the beam-beam interaction has been studied for both beams. As expected, due to the limited bunch current of the electrons, the proton beam-beam effect has only a very marginal impact on the proton emittance and stability. In the end, the electron bunch intensity is in the order of one percent of the proton bunch population, which leads to a very limited beam-beam force compared to the other LHC proton-proton interaction points. The linear part of the beam-beam interaction as well as the non-linear effect has been studied and long-range forces are not present: due to the fast beam separation scheme, the beam separation between electrons and protons is already well beyond 20 $\sigma$ beam size, and thus – based on LHC operation experience – long-range effects can be neglected. Studies on long-term stability, including orbit fluctuations of the electron beam, are ongoing and might support the need for a fast acting correction system.
Beam-beam effect on the proton beam. The beam-beam interaction on the other side has a highly unbalanced effect, as it depends on the energy and the bunch intensity of the opposing beam. In linear approximation the resulting tune-shift is given by:
@@
where we refer here to the effect on the proton beam; accordingly $N_{e}$ is the electron bunch population, $r_{0}$ the classical proton radius, $\gamma$ is the electron Lorentz factor and $\beta^{\ast}$ and $\sigma_{x,y}$ are the amplitude function and transverse beam sizes at the Interaction Point (IP). We would like to point out that in this approximation the tune-shift does not depend on the $\beta$-function. However, it depends on the energy of the beam considered. For the design case of the 50 GeV electron and 7 TeV proton beams, we estimate the effect using a realistic MAD-X modelling of the HL-LHC lattice including the electron quadrupoles in the IR, and the remaining distortions are compensated; see Fig. 4. Due to the low energy of the electrons – compared to the 7 TeV of the proton beam – the insertion quadrupoles of the electron lattice have only a marginal effect on the proton beam: the resulting beam-beam tune-shift of $\Delta Q_{x,y}\approx 3.8\cdot 10^{-4}$ is nearly two orders of magnitude smaller than the effect of the proton-proton interactions for the HL-LHC operation where we obtain $\Delta Q_{x,y}\approx-8.6\cdot 10^{-3}$ per IP and thus the additional distortion due to $ep$ collisions only has a marginal effect. The distortion of the proton beta function obtained is only $1.8\%$ – without correction. It is corrected nevertheless by slightly modifying the gradients of the neighbouring p-quads (a so-called local rematch of the beam optics).
@@
Beam-beam effect on the electron beam. The situation changes considerably for the electrons. Due to the lower beam energy the electrons are more sensitive, and especially the effect of the non-linear terms has to be minimised. In Fig. 5, we compute the beam size of the electron beam as a function of the position $s$ along the beam orbit with $s=0$ referring to the IP. In black, the theoretical beam size is shown as it is obtained without beam-beam forces. The strong additional focusing due to the space-charge forces of the proton bunch is also shown in blue. The green dashed line quantifies the mismatch with the design linear optics [10]. Here we would like to point out that the dominant parameter, the proton bunch population, refers to the ultimate value of $N_{p}=2.2\cdot 10^{11}$ as foreseen for the HL-LHC luminosity upgrade. Under the influence of this additional focusing – called “pinch effect” – the electron beam reaches a smaller effective beam size which is, however, to a large extent compensated by the so-called hourglass effect in the vicinity of the IR. However, the optics distortion has to be corrected in order to obtain well-matched beam parameters before entering the return arc and the decelerating process of the ERL. This is also highlighted in Fig. 5 where, using the same modelling of the modified HL-LHC lattice introduced above, the black and blue curves are shown to be congruent before and after the interaction region.
@@
Figure 5: Effect of the space charge force on the 20 GeV electron beam, or “beam-beam force”. The beam size is reduced (leading to an additional enhancement of the luminosity, pinch effect) and re-matched to the ideal conditions before entering the return arc for energy recovery.
@@
Electron emittance. The above mentioned re-optimisation of the beam-beam interaction clearly involves matched beam sizes between electron and proton beam in both planes:
@@
Thus, the emittance of the electrons for the 20 GeV case has to be taken into account in a proper manner. In a storage ring the equilibrium emittance is well defined by the emission of synchrotron light. In the case of an ERL however, this equilibrium is never reached. Ignoring the small effect of quantum emission in return arcs, the electron emittance therefore follows the rule of adiabatic damping $\varepsilon\propto 1/\gamma$. As a consequence, the emittance of the 20 GeV electron beam at the interaction point is considerably larger than the corresponding value at 50 GeV. This fact has to be taken into account to fulfil the matching condition for the transverse sizes between protons and electrons at the IP. In principle a strong reduction of the $\beta$-function would be needed to compensate for the increased beam emittance. However, this is far beyond the capabilities of the mini-beta quadrupoles.
In addition, the conditions of matched beam sizes for the 20 GeV phase-one scenario are obtained with electron beam optics we slightly adjusted from the final LHeC configuration for a $\beta^{*}$ of 20 cm. Given the emittance and amplitude function of the colliding proton beam that define the optics conditions at the IP, we get a beam size at the $ep$ interaction point of:
@@
and the electron beam has to follow this condition.
For the amplitude function of the electrons, $\beta^{\ast}=20$ cm we obtain a required absolute (i.e. geometric) electron emittance of $\varepsilon_{0}=5.7\cdot 10^{-4}~{\rm mm\cdot mrad}$. This corresponds to a normalised (energy-independent) value of $\varepsilon_{n}=22~{\rm mm\cdot mrad}$ and lies well within the specifications of the proposed particle source for the ERL [2]. According to the design parameters of the electron gun a normalised emittance even as low as $\varepsilon_{n}=6~{\rm mm}\cdot{\rm mrad}$ is expected (see Tab. 1) which represents a comfortable safety margin for the 20 GeV design.
Table 1: General specification of the LHeC ERL electron source.
@@
In these considerations, the small effect of the emittance increase in the return arcs of the ERL was taken into account. Given the lower beam energy of 20 GeV, the emitted quanta have a much smaller effect, and the resulting emittance dilution remains at a very low level of a few percent. Other effects might influence the quality of the electron beam: the effect of CSR (coherent synchrotron radiation) has been studied in the context of front-to-end simulations that have been performed for the ERL. Given the parameter choice of the machine no significant influence has been found. Beam-Breakup (BBU) as well as micro bunching instabilities on the other side are items still on the to-do-list towards a technical design report as well as tolerance considerations for the beam-beam effect between the non-colliding proton beam and the electrons, for example. They will be investigated in the near future.
Beam separation scheme and synchrotron radiation. The beam separation between electrons and protons is provided by the combined effect of the detector weak dipole field and the off-centred quadrupoles of the electron mini-beta concept. Marked in grey in Fig. 3 the combined function quadrupoles of the doublet and the dipole field of the detector represent a bending force with equal and constant bending radius. They provide a smooth separation effect because of the different beam rigidities of electron and proton beams with a minimised critical energy and limited power of the emitted synchrotron light. Applying this scheme to the case of 20 GeV electron energy results in $E_{\rm crit}\approx 15-20~{\rm keV}$ and an overall radiated power in the interaction region of well below 1 kW. Optimisation work is performed modelling the electron lattice including the first HL-LHC proton quadrupoles close to the interaction region, and using version 1.7.7 of the beam delivery simulation toolkit (BDSIM) [11], With a reduced region of interest centred on the first elements around IP2 (including the first proton quadrupoles), it assumes two scenarios for the free drift length: $L^{\ast}=21$, and $15~{\rm m}$. Although the reduction of the radiated power is clear between these two values of the $L^{\ast}$ for the 50 GeV baseline, it is observed to be negligible for the 20 GeV option.
@@
Design, work load, and cost optimisation. Following the idea of a design that can be easily extended in energy by staging additional return arcs, the basic layout of the ERL will remain the same. Starting with one-pass ERL the return arcs from 3 to 6 as well as the so-called beam spreaders and re-combiners at the end of the linacs and the bypass that guides the low-energy beam outside the particle detector will not be needed. An estimate of the material cost for the original ERL design resulted in a magnet and vacuum system budget of $11~{\rm kCHF}$ per metre [12]. As a result, an overall budget reduction of at least $70~{\rm MCHF}$ is obtained for the 20 GeV version discussed here. In addition, the work load and associated personnel costs are significantly reduced. Moreover, the lower electron energy allows a much simplified design of the interaction region – in the baseline case, with $\rho\approx 1500~{\rm m}$, the synchrotron radiation is relatively weak and does not require installation of sophisticated protection masks and absorbers. Finally, as discussed in the following, detector costs are minimised by joining forces with the ALICE 3 project. Last but not least, the operating costs of the 20 GeV ERL will be much smaller with respect to the 50 GeV case because of much lower energy losses due to synchrotron radiation and the absence of the arc 3–6 beamlines.
As discussed in the 2021 CDR revision of the LHeC study [2] the installation of the ERL machine can be made partially in parallel during the Run4 operation period and the detector installation would then take place during the fourth Long Shutdown (LS4) planned in 2035–6 [13]. In total, 7 years are needed to build the LHeC, including permits, engineering studies, tendering and land negotiations, and the final year of connection to HL-LHC [4]. Therefore, given a much simpler design of the phase-one LHeC proposed here and a timely decision on the start of such a project, we believe that the electron-hadron collisions can be available during Run5 of the HL-LHC which will start, according to the present planning, in 2036.
@@
Polarisation. Polarised electron beams can be rather naturally incorporated into an ERL type accelerator. Laser-driven state-of-the-art electron guns provide polarised beams up to 90% as previously suggested in the LHeC CDR [2]). Harmful effects like resonant depolarisation that play a major role in storage rings are by definition not existent, and the de-tuning of the spin precession due to the emitted synchrotron light is estimated to be on the order of a few percent. As a consequence the acceleration chain provided by the ERL is quasi-spin transparent and does not play a large role in the final – longitudinal – electron beam polarisation at the IP.

### v3

@@ 1 Introduction
The Large Hadron-electron Collider (LHeC) [LHeCStudyGroup:2012zhm, LHeC:2020van] will open new frontiers in high-energy physics as it will become a unique laboratory for studies of deep-inelastic scattering (DIS) and Higgs boson, electroweak, top quark and beyond the Standard Model (BSM) phenomena in a vast range of virtualities $Q^{2}$, from very close to zero up to $10^{6}$ GeV2, and for a much extended range of $x$ – the fractional longitudinal momentum of partons. Recently, an intense electron beam from a novel energy recovery linac (ERL) was proposed to collide with a proton or ion beam from the circular, 27-km circumference High-Luminosity Large Hadron Collider (HL-LHC) at CERN, in the redesigned P2 interaction region capable of hosting electron-hadron or hadron-hadron collisions, concurrently with hadron-hadron collisions in other HL-LHC experiments [Andre:2022xeh].
In the Conceptual Design Report (CDR) of the LHeC, it is assumed that a three-pass ERL delivers a 50 GeV electron beam with an intensity of up to 20 mA [LHeC:2020van], and in the most recent update of the LHeC proposal, its commissioning is proposed several years after the completion of the HL-LHC programme [Ahmadova:2025vzd]. In this paper, we propose the use of a 20 GeV single-pass ERL with a higher beam current of 60 mA, and with a circumference of a third of the LHC’s. This choice of design is motivated by the availability of the technology to operate a single-pass ERL (although presently at lower beam currents) and by the overall simplification, in particular, in the design of the adequate machine-detector interface owing to much lower levels of synchrotron radiation, allowing much earlier commissioning of the high energy electron-hadron collisions at CERN. The electron beam will be accelerated to a full energy of 20 GeV by one pass in two 10 GeV linacs (see Fig. 1), and after collision with the clockwise circulating LHC proton beam, it will be guided back to the linac structures in counter-phase to be completely decelerated within the energy recovery process.
@@
The technical feasibility of such a single-pass design has been demonstrated in a number of test facilities and linacs that operate in the range of several mA beam current (see, e.g., [Hutton:2022kac]). Fig. 2 summarises ERLs that are already operational, in the phase of commissioning or planned.
Figure 2: Different ERL projects worldwide that are already operational or in their commissioning / planning phase. The red and green markers indicate the baseline LHeC design and the ERL configuration described in this paper, respectively. Figure adapted from [Hutton:2022kac].
Note that the cost of building this ”phase-one” LHeC as well as its running costs are significantly lower than for the final LHeC configuration. Such a staged realisation of the LHeC project provides a huge benefit not only in reducing financial (and personnel) strain but first of all in the availability of concurrent electron-hadron collisions already during Run5 of the current HL-LHC programme. Consequently, apart from the possibility of carrying out a wide range of unique and important studies almost 10 years earlier, the results of this electron-hadron experiment could also be used to improve many important measurements in other (hadron-hadron) experiments at the HL-LHC in a timely manner.
2 The LHeC IR
@@
Schematically, the concept is shown in Fig. 1 – the geometry of the ERL will remain the same: two superconducting linacs are connected by return arcs on both sides – details about the general layout can be found in [Holzer:2022eac] and [Bogacz:2022ufs]. However, for the phase-one case with reduced electron energy several parts needed to reach higher energies are left out: arcs from 3 to 6 will not be installed, nor will the so-called bypass needed in the final 50 GeV design to guide the beam outside the particle detector, which will only be installed in a later staging step. Beam spreaders and re-combiner modules will also not be present in the phase-one electron beam lattice.
On the other hand, a number of key elements will remain unchanged for both ERL phases. The linacs, the interaction region (IR) and the beam separation scheme will be designed to be valid for any operation energy. As described in detail in [Holzer:2022eac, vonWitzleben:2023yca] the proton and electron beams are separated after the interaction point through their different beam rigidities. A combination of a weak dipole field embedded inside the detector and the combined function effect of the off-centred mini-beta quadrupoles of the electron lattice acts as an effective separation scheme. The actual value of the dipole strength and the offset of the quadrupoles were optimised for the lowest synchrotron radiation and thus equal effective bending radii in the three magnets [Andre:2022zqx], resulting in a bending radius of the electron trajectory in this separation scheme of about 1500 m.
@@
Figure 3: The interaction region of electron and proton beams: in black the mini-beta focusing scheme of the electrons is shown, including the detector dipole field that is part of the beam separation. Both are embedded within the free space of the HL-LHC proton inner triplet lattice, marked in blue and red. Figure adapted from [tizi].
The interaction region of the $ep$ collisions is shown schematically in Fig. 3. Given the comparatively low energy of the electrons, a gradient of approximately g=20g=20~T/m and a length of 1.8 m are sufficient to meet the optics requirements of the electron quadrupoles. These magnets can therefore be made very compact, and the dimensions of the electron mini-beta insertions have been chosen so that they will fit easily in the free space of the IR. Only the first quadrupole of the proton lattice, “Q1” in the figure, will have to be replaced by a new design that allows a field free region for the electrons as their separation is not large enough yet to be guided completely outside the cryostat. For details of the technical design of this new magnet we refer to [LHeC:2020van]. With the separation scheme explained here, head-on collisions between the proton and electron beam are foreseen, to avoid luminosity reduction because of the geometric loss factor from an eventual crossing angle.
@@
Figure 4: Local distortion of the proton optics due to the influence of the electron quadrupoles before (red) and after a local compensation of the effect (blue) for the case of a 20 GeV electron energy. Figure extracted from [tizi].
Beam-beam effect on the proton beam. Beyond the optics distortion from the electron gradient fields that have a certain influence on the proton optics, the effect of the beam-beam interaction has been studied for both beams. The beam-beam interaction has a highly unbalanced effect, as it depends on the energy and the bunch intensity of the opposing beam. In linear approximation the resulting tune-shift for the proton beam is given by:
@@
where $N_{e}$ is the electron bunch population, $r_{0}$ the classical proton radius, $\gamma$ is the electron Lorentz factor and $\beta^{\ast}$ and $\sigma_{x,y}$ are the amplitude function of the protons and the transverse beam sizes at the Interaction Point (IP). We would like to point out that in this approximation the tune-shift does not depend on the $\beta$-function. However, it depends on the energy of the beam considered. As expected, due to the limited bunch current of the electrons, the proton beam-beam effect has only a very small impact on the proton emittance and stability. In the end, the electron bunch intensity is in the order of one percent of the proton bunch population, which leads to a very limited beam-beam force compared to the other LHC proton-proton interaction points. The linear part of the beam-beam interaction as well as the non-linear effect has been studied and long-range forces are not present: due to the fast beam separation scheme, the beam separation between electrons and protons is already well beyond 20 $\sigma$ beam size at the first parasitic bunch crossing and thus – based on LHC operation experience – long-range effects can be neglected. Studies on long-term stability, including orbit fluctuations of the electron beam, are ongoing and might support the need for a fast acting correction system.
Given the parameters of the LHeC, the resulting beam-beam tune-shift of $\Delta Q_{x,y}\approx 3.8\cdot 10^{-4}$ is nearly two orders of magnitude smaller than the effect of the proton-proton interactions for the HL-LHC operation where we obtain $\Delta Q_{x,y}\approx-8.6\cdot 10^{-3}$ per IP and thus the additional distortion due to $ep$ collisions only has a marginal effect.
@@
Beam-beam effect on the electron beam. The situation changes considerably for the electrons. Due to the lower beam energy the electrons are more sensitive, and especially the effect of the non-linear terms has to be minimised. In Fig. 5, we compute the beam size of the electron beam as a function of the position $s$ along the beam orbit with $s=0$ referring to the IP. In black, the theoretical beam size is shown as it is obtained without beam-beam forces. The strong additional focusing due to the space-charge forces of the proton bunch is also shown in blue. The green dashed line quantifies the mismatch with the design linear optics [Hao:2014yoa]. Here we would like to point out that the dominant parameter, the proton bunch population, refers to the ultimate value of $N_{p}=2.2\cdot 10^{11}$ as foreseen for the HL-LHC luminosity upgrade. Under the influence of this additional focusing – called “pinch effect” – the electron beam reaches a smaller effective beam size which is, however, to a large extent compensated by the so-called hourglass effect in the vicinity of the IP. However, the optics distortion has to be corrected in order to obtain well-matched beam parameters before entering the return arc and the decelerating process of the ERL. This is also highlighted in Fig. 5 where, using the same modelling of the modified HL-LHC lattice introduced above, the black and blue curves are shown to be congruent before and after the interaction.
@@
Figure 5: Effect of the space charge force on the 20 GeV electron beam, or “beam-beam force”. The beam size is reduced (leading to an additional enhancement of the luminosity, pinch effect) and re-matched to the ideal conditions before entering the return arc for energy recovery. The best re-matching was obtained for non-zero value of the slope of amplitude function at the IP, $\alpha^{*}=-0.2$.
@@
Electron transverse emittance. The above mentioned re-optimisation of the beam-beam interaction involves matched beam sizes between electron and proton beam in both planes:
@@
Thus, the transverse emittance of the electrons for the 20 GeV case has to be properly considered. In a storage ring the equilibrium emittance is well defined by the emission of synchrotron light. In the case of an ERL however, this equilibrium is never reached. Ignoring the small effect of quantum emission in return arcs, the electron transverse emittance therefore follows the rule of adiabatic damping $\varepsilon\propto 1/\gamma$. As a consequence, the emittance of the 20 GeV electron beam at the interaction point is considerably larger than the corresponding value at 50 GeV. This fact has to be taken into account to fulfil the matching condition for the transverse sizes between protons and electrons at the IP. In principle a strong reduction of the $\beta$-function would be needed to compensate for the increased beam emittance. However, this is far beyond the capabilities of the mini-beta quadrupoles and the matching of electron beam-size has to be achieved by using the electron source with a reduced transverse emittance.
In addition, the conditions of the matched beam sizes for the 20 GeV phase-one scenario are obtained with the proton beam optics for the LHC Run5, with a $\beta_{p}^{*}$ of $35~{\rm cm}$. Given the emittance and amplitude function of the colliding proton beam, one gets a beam size at the $ep$ interaction point:
@@
which is matched by the electron beam, assuming an absolute (i.e. geometric) electron emittance of $\varepsilon_{0}=5.7\cdot 10^{-4}~{\rm mm\cdot mrad}$ at the energy of 20 GeV which corresponds to a normalised (energy-independent) value of $\varepsilon_{n}=22~{\rm mm\cdot mrad}$ and lies well within the specifications of the proposed particle source for the ERL, see Tab. 1 . In order to fulfil the requirements of matched beam sizes, a beam optics with $\beta_{e}^{\ast}=20$ cm has been established that includes – as explained above – the beam-beam effect, namely the additional focusing due to the space charge of the proton bunches, see Fig. 5. We would like to emphasise in this context that, unlike the proton optics of the HL-LHC where a manifold of boundary conditions had to be taken into account, the optics of the final focus system of the ERL is much more flexible. Consequently, as long as an optics rematch to the return arc 2 is provided to prepare the electron beam for the decelerating phase, a variety of beam optics can be considered.
Table 1: General specification of the LHeC ERL electron source [LHeC:2020van].
@@
In these considerations, the small effect of the emittance increase in the return arcs of the ERL was taken into account. Given the lower beam energy of 20 GeV, the emitted quanta have a much smaller effect, and the resulting emittance dilution remains at a very low level of a few percent. Other effects might influence the quality of the electron beam: the effect of CSR (coherent synchrotron radiation) has been studied in the context of front-to-end simulations that have been performed for the ERL. Given the parameter choice of the machine no significant influence has been found.
Beam separation scheme and synchrotron radiation. The beam separation between electrons and protons is provided by the combined effect of the detector weak dipole field and the off-centred quadrupoles of the electron mini-beta concept. Marked in grey in Fig. 3 the combined function quadrupoles of the doublet and the dipole field of the detector represent a bending force with equal and constant bending radius. They provide a smooth separation effect because of the different beam rigidities of electron and proton beams with a minimised critical energy and limited power of the emitted synchrotron light. Applying this scheme to the case of 20 GeV electron energy results in $E_{\rm crit}\approx 15-20~{\rm keV}$ and an overall radiated power in the interaction region of well below 1 kW. Optimisation work is performed modelling the electron lattice including the first HL-LHC proton quadrupoles close to the interaction region, and using version 1.7.7 of the beam delivery simulation toolkit (BDSIM) [Nevay:2018zhp], As for the 50 GeV baseline, two free drift lengths $L^{\ast}=21$, and $15~{\rm m}$ scenarios are studied in this 20 GeV option for the reduced region of interest centred on the first elements around IP2 (including the first proton quadrupoles), highlighted in grey in Fig. 3. Although the reduction of the radiated power is clear between these two values for the 50 GeV baseline, it is observed to be negligible for the 20 GeV option.
@@
Design, work load, and cost optimisation. Following the idea of a design that can be easily extended in energy by staging additional return arcs, the basic layout of the ERL will remain the same. Starting with one-pass ERL the return arcs from 3 to 6 as well as the so-called beam spreaders and re-combiners at the end of the linacs and the bypass that guides the low-energy beam outside the particle detector will not be needed. An estimate of the material cost for the original ERL design resulted in a magnet and vacuum system budget of $11~{\rm kCHF}$ per metre [Bruning:2652349]. As a result, an overall budget reduction of at least $70~{\rm MCHF}$ is obtained for the 20 GeV version discussed here. In addition, the work load and associated personnel costs are significantly reduced. Moreover, the lower electron energy allows a much simplified design of the interaction region – in the baseline case, with $\rho\approx 1500~{\rm m}$, the synchrotron radiation is relatively weak and does not require installation of sophisticated protection masks and absorbers. Finally, as discussed in the following, detector costs are minimised by joining forces with the ALICE 3 project. Last but not least, the operating costs of the 20 GeV ERL will be much smaller with respect to the 50 GeV case because of much lower energy losses due to synchrotron radiation and the absence of the arc 3–6 beamlines.
As discussed in the 2021 CDR revision of the LHeC study [LHeC:2020van] the installation of the ERL machine can be made partially in parallel during the Run4 operation period and the detector installation would then take place during the fourth Long Shutdown (LS4) planned in 2035–6 [lhc_run_planning]. In total, 7 years are needed to build the LHeC, including permits, engineering studies, tendering and land negotiations, and the final year of connection to HL-LHC [Ahmadova:2025vzd]. Therefore, given a much simpler design of the phase-one LHeC proposed here and a timely decision on the start of such a project, we believe that the electron-hadron collisions can be available during Run5 of the HL-LHC which will start, according to the present planning, in 2036.
@@
Polarisation. Polarised electron beams can be rather naturally incorporated into an ERL type accelerator. Laser-driven state-of-the-art electron guns provide polarised beams up to 90% as previously suggested in the LHeC CDR [LHeC:2020van]. Harmful effects like resonant depolarisation that play a major role in storage rings are by definition not existent, and the de-tuning of the spin precession due to the emitted synchrotron light is estimated to be on the order of a few percent. For details see [Bai:1381541]. As a consequence the acceleration chain provided by the ERL is quasi-spin transparent and does not play a large role in the final – longitudinal – electron beam polarisation at the IP.